OpenStudy (anonymous):

Evaluate the following function at the values x+1 and x+h and find f(x+h)-f(x)/h

OpenStudy (anonymous):

so... f(x+1)=

OpenStudy (anonymous):

oops, forgot to mention the main equation of f(x)=xsquared+5

OpenStudy (anonymous):

Plug x+1 into the function f(x) = x^2 + 5 that is take the x + 1 and put it where the x is. f(x+1) = (x+1)^2 +5 Foil the (x+1)^2. Add 5 to simplify. Now do the same with the x+h.

OpenStudy (anonymous):

so f(x+1)=xsquared-5x-5?

OpenStudy (anonymous):

Not quite, but almost! (x+1)(x+1)-5 = x^2 +2x +1 - 5 which equals what?

OpenStudy (anonymous):

\[x ^{2}+4x-4?\]

OpenStudy (anonymous):

Good! now you're going to plug in x + h just like you did the x + 1.

OpenStudy (anonymous):

okay, thank you!

OpenStudy (anonymous):

Did you get x^2 + 2xh + h^2 - 5 ?

OpenStudy (anonymous):

Finally the last bit. [f(x+h) - f(x)]/h = [x^2 + 2xh + h^2 -5 - (x^2 - 5)]/h Run the negative through the x^2 - 5 then simplify the numerator. You'll find that when you factor out the common h then you can cancel the h in the denominator. This is the precursor to limits in calculus so make sure you understand this concept thoroughly before continuing!